OpenStudy (anonymous):

Solve square root x-1=3

OpenStudy (anonymous):

\[\sqrt{x-1}=3\]

OpenStudy (anonymous):

\[\sqrt{x-1}^{2}=3^{2}\] x-1=9 +1 +1 x=10